### v1

# pnasresearcharticle

Abstract
Hydrogels are crosslinked polymer networks with high water content widely utilized in biomedical applications. Injectable depot-forming hydrogels for drug delivery offer sustained release of biomolecular cargo, leveraging controlled macromolecular diffusion within dynamic networks to promote extended immunity and favorable immunological responses. However, mitigating burst release while maintaining sustained-release properties remains a challenge due to network disruption during needle extrusion, leading to excessive short-term cargo release and local toxicity. While controlling both short- and long-term release profiles would mitigate this issue, it remains a persistent challenge. Here, we present a hydrogel formulation strategy for self-regulating, post-injection network evolution enabled by exchange of functional molecules with physiological environment, achieving controlled cargo release over both short and long durations. By tuning network viscoelasticity, we optimized rheological properties to minimize burst release while preserving sustained-release profiles. Using model protein cargo, in vitro and in vivo studies demonstrated reduced burst release and comparable sustained-release to unmodified control formulations. These findings highlight a novel strategy for precise cargo delivery in biomedical applications.
keywords:
Hydrogels |||| Cargo diffusion |||| Drug delivery |||| Rheology
[…]
This manuscript was compiled on March 19, 2025 www.pnas.org/cgi/doi/10.1073/pnas.XXXXXXXXXX
7]3
[…]
Hydrogels are crosslinked polymer networks with high water content, known for their biocompatibility and in vivo tolerability. Their highly tunable mechanochemical properties make them versatile materials for diverse applications (1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12). In the biomedical field, hydrogels have garnered significant attention, leading to discoveries of novel properties and applications (1, 3, 5, 6, 7, 9, 10, 11). Injectable, depot-forming hydrogels represent a critical subfield, enabling minimally invasive, sustained drug delivery. Using the slower diffusion of macromolecules, including proteins and other biologics, from dynamic hydrogel networks, these systems achieve prolonged therapeutic release, often far exceeding the durations of intravenous administration (cf. Fig. 1(a)), and elicit favorable immunological responses compared to bolus injections, underscoring their potential in advanced therapeutic applications (7, 13, 14, 15, 9, 6, 10, 3, 4, 16, 11).
A significant body of research focuses on developing hydrogels that prolong cargo retention in injected depots by tuning rheological properties (stiffness, relaxation time, yielding, and creep behavior) that control injection, depot persistence, and mass transport. The goal is to limit erosive cargo release and engineer a slow diffusive release rate (3, 13, 10, 17, 18, 19, 7). A typical sustained release profile from injectable hydrogel depots is shown in Fig. 1(a) (red). Physically crosslinked hydrogels with shear-thinning and self-healing properties enable easy injection and depot formation. Once injected, the release rates and cargo retention depend on the network structure, governed by its linear viscoelastic rheological properties (2, 3, 4, 20, 21, 5, 22, 23, 24, 16, 11). Increasing the stiffness of hydrogel networks (plateau shear modulus $G_{0}$) decreases the cargo release rate by reducing the effective mesh size ($\xi$), which impedes solute diffusion and lowers the diffusion coefficient compared to its Stokes-Einstein diffusivity ($\mathbb{D}/\mathbb{D}_{0}$). This relationship is supported by naïve network theory, which predicts $\xi\sim G_{0}^{-1/3}$, a trend confirmed by more advanced models (25, 26, 20, 19, 20, 13, 17, 23, 26, 16, 19).
The focus has primarily been on prolonging the duration of release by retaining cargo within the depot, rather than tailoring release profiles to control both short-term burst release and sustained release over longer periods (7, 6, 3, 10, 16, 1, 11). Despite excellent sustained release characteristics, most injectable hydrogel platforms exhibit burst release, leading to rapid absorption and therapeutically harmful local concentrations of drugs or antibodies, and high peak plasma concentrations, which may cause adverse effects (27, 28). Furthermore, burst release during the first few hours post-injection, as the hydrogel recovers its stiffness, reduces the duration over which the release remains within physiologically effective levels. The key challenge lies in minimizing burst release while maintaining appropriate sustained release profiles. Decoupling these two temporal regimes of mass transport to separately engineer each release phase presents significant rheological challenges.
Figure 1: (a) Primary objective: modulate cargo release by designing the structural evolution of injectable hydrogels, minimizing burst release and extending sustained release. (b) crosslinking dynamics transition from fast to slow, altering cargo diffusion rates. Fast-relaxing networks reduce burst release, while slow-relaxing networks decrease diffusion coefficients, enhancing prolonged retention and sustained release. (c) Chemical structures and microstructures of the hydrogels: transiently crosslinked PVA-Borax networks formed via covalent associative bonds between borate ions and PVA groups, and Sangelose hydrogels formed by hydrophobic clustering of stearyl groups on HPMC chains into dynamic, noncovalently crosslinked networks.
Our observations indicate that the dynamics of hydrogel networks, specifically the lifetimes and kinetics of the dynamic crosslinks that govern the self-diffusion of the network, play a key role in cargo diffusion, beyond the influence of topology (e.g., the average mesh size related to stiffness) (cf. Fig. 2). Although this phenomenon has been observed in previous studies (23, 24, 22), it remains underexplored. We propose that hydrogel structure recovery post-injection, during which crosslinks dissociate due to large strains in the narrow flow geometry, correlates with relaxation behavior: networks that relax stress rapidly also recover their structural integrity faster by reforming crosslinks. Hydrogels with slow crosslink kinetics exhibit favorable sustained release characteristics as a result of reduced cargo diffusion from their fully recovered structure. However, their slower recovery post-injection leads to excessive burst release, as loosely connected networks fail to arrest rapid diffusive or advective solute release driven by steep concentration gradients in cargo-free environments, such as intramuscular or subcutaneous spaces. In contrast, hydrogels with fast crosslink kinetics recover structure quickly, mitigating burst release by resisting mass transport. However, these materials often exhibit higher diffusion coefficients, compromising sustained release compared to slower relaxing networks.
Figure 2: Transport properties of the hydrogels used. Frequency sweep linear viscoelastic rheology data for (a) PVA-Borax and (b) Sangelose hydrogels. First plot in each row: experimental and fitted linear viscoelastic moduli ($G^{\prime}$: filled circles, solid line; $G^{\prime\prime}$: open circles, dashed line) using log-normal (PVA-Borax) or modified BSW (Sangelose) relaxation spectra. Second plot: corresponding spectra with total (solid), viscoelastic/flow (dashed), and glassy (dotted) components. Bar plots show stiffness ($G_{0}$) and relaxation time ($\tau_{\rm R}$), highlighting hydrogel rheological tunability. (c) Cargo diffusivity measured by fluorescence recovery after photobleaching. Grayscale images show photobleaching and recovery, with exponential fits providing diffusion coefficients ($\mathbb{D}$) normalized to the free-solvent value ($\mathbb{D}_{0}$). Data reveals correlations between diffusivity, stiffness, and relaxation dynamics.
Achieving control over both short- and long-term cargo release profiles is thus challenging, often resulting in a compromise between minimizing burst release and ensuring prolonged release. In this work, we propose a strategy to address this challenge, as shown in Fig. 1(b). To reduce burst release, networks with fast crosslink kinetics are used initially for the injections. Subsequently, the network is allowed to evolve so the kinetics are slowed to achieve sustained release profiles similar to those of kinetically slow networks. The ideal release profile is shown in Fig. 1(a) (blue). Therefore, designing a hydrogel system that evolves its structure post-injection, transitioning from fast to slow crosslinking kinetics throughout the release period, is crucial.
Results and Discussion
Figure 3: In vitro cargo release studies. (a) Experimental setup: cargo-loaded hydrogel injected into buffer, with buffer collected over time to quantify protein release. (b) Hypothesized mechanisms for burst and sustained release in PVAB1/buffer1 and PVAB3/buffer3 systems. (c) Cumulative release profiles ($M_{t}/M_{0}$) for PVA-Borax hydrogels with BSA cargo in respective buffers, with inset showing linear-scale data. (d) Relative areas under burst and sustained release curves for PVA-Borax. (e) Cumulative release profiles ($M_{t}/M_{0}$) for Sangelose hydrogels with hIgG cargo in PBS, with linear-scale inset. (f) Relative areas under burst and sustained release curves for Sangelose.
Figure 4: In vitro cargo release studies with network evolution. (a) Experimental setup: cargo-loaded hydrogel injected into buffer, with buffer collected over time to quantify protein release. (b) Hypothesized mechanisms of burst and sustained release for PVAB1/buffer3 and PVAB3/buffer1 systems, illustrating network evolution leading to (i) favorable or (ii) unfavorable release. (c) Cumulative release profiles ($M_{t}/M_{0}$) for PVA-Borax hydrogels with BSA cargo in opposite buffers, with linear-scale inset. (d) Relative areas under burst and sustained release curves for PVA-Borax.
Hydrogels with Modular Transport Properties: Linear Rheology and Diffusivity
The timescale of network relaxation, directly related to crosslinking kinetics, provides an additional degree of freedom in tuning cargo diffusivity within hydrogel meshes (22, 29). For networks with similar stiffness, those with distinct stress relaxation times exhibit different solute diffusivity, as observed in the hydrogels used in this study (cf. Fig. 2). This property is leveraged to design materials with rheological properties that evolve between fast and slow crosslink kinetics, controlling solute diffusion while maintaining similar stiffness. To achieve this, small molecules, such as specific ions or surfactants, are incorporated to modulate crosslinking and promote faster network relaxation. While this may enhance burst release, a slow diffusion of these molecules from the network allows the structure to gradually regain slower crosslinking kinetics, lowering the diffusion coefficient and prolonging cargo retention for sustained release.
To demonstrate that networks with similar stiffness but different relaxation times exhibit varying cargo diffusivity (Fig. 2(a, c)), we formulated a model transient hydrogel network (Fig. 1(c)). The system consisted of poly(vinyl alcohol) chains covalently crosslinked by borate ions. Borate- and boronic ester-crosslinked hydrogels are associative networks, where crosslinking kinetics directly influence relaxation behavior (30, 23, 31). The relaxation time and dissociation rate constant are related as $\tau_{\rm R}\simeq k_{\rm d}^{-1}$, as stress relaxation occurs via crosslink dissociation and reformation. Slowing dissociation kinetics significantly delays network relaxation. Modulating $p$H above the boric acid $pK_{\rm a}\simeq 9$ alters the crosslink dissociation kinetics, slowing relaxation without changing network stiffness, which remains governed by the average molecular weight between crosslinks. This is due to the $p$H-induced shift in association ($k_{\rm a}$) and dissociation ($k_{\rm d}$) rate constants, maintaining a constant equilibrium constant $K_{\rm eq}=k_{\rm a}/k_{\rm d}$, which is correlated with $G_{0}$. This orthogonal control over stiffness and relaxation time allows us to isolate the effect of network relaxation on cargo diffusion and formulate materials with comparable stiffness but different relaxation dynamics (cf. Fig. 2(a)).
The effect of network topology and dynamics on cargo release from hydrogels was bench-marked using in vitro release assays (13). We injected two PVA-Borax hydrogels with different relaxation times but similar stiffness ($G_{0}\simeq 675\leavevmode\nobreak\ {\rm Pa}$ and $735\leavevmode\nobreak\ {\rm Pa}$), into capillary tubes with buffers of matching pH and hydroxyl ion concentrations. This setup ensures a dynamic equilibrium between the hydrogel depot and the buffer interface, allowing only slow diffusive release of protein cargo. Cumulative release data ($M_{t}$) over 14 days, relative to the total cargo ($M_{0}$), is shown in Fig.3(c). To quantify these differences, we calculated the relative areas under the release profiles for both burst and sustained regimes (cf. Fig. 3(d)), such that $\alpha_{\rm burst}/\alpha_{\rm tot}$ and $\alpha_{\rm sust}/\alpha_{\rm tot}$, where $\alpha_{T_{1}T_{2}}=\sum_{n=T_{1}}^{T_{2}}(M_{t})_{n}t_{n}$ is the area between times $T_{1}$ and $T_{2}$, and $\alpha_{\rm tot}=\sum_{n=1}^{\rm end}(M_{t})_{n}t_{n}$ is the total area. For burst release, $T_{1}=0$ and $T_{2}=1\leavevmode\nobreak\ {\rm day}$; for sustained release, $T_{1}=2\leavevmode\nobreak\ {\rm days}$ and $T_{2}=14\leavevmode\nobreak\ {\rm days}$. As hypothesized, PVAB1/buffer1 releases less cargo in the burst phase due to its faster network relaxation, but more cargo over longer durations compared to PVAB3/buffer3. If PVAB1/buffer1 were functionalized to evolve into a slower relaxing network, we could potentially achieve both suppressed burst release and improved sustained release (vide infra).
In Vitro Cargo Release of Evolving Hydrogels
[…]
Figure 5: In vivo pharmacokinetics studies with Sangelose hydrogels containing hIgG cargo. (a) BL6 mice injected with hIgG-loaded hydrogels, with serum protein levels analyzed at specific time points. (b) Hypothesized release mechanism: structure evolution in S3Tw2 (with Tween 80) enables favorable short- and long-term release. (c) Serum protein concentration profiles for Sangelose hydrogels, showing burst release and sustained phases; solid lines represent two-compartment PK model fits. (d) Cumulative release across burst and sustained phases, with S3Tw2 demonstrating superior performance. (e) Cargo release half-lives, with S3Tw2 achieving slower release due to improved sustained release and reduced burst effects compared to S3Tw0.
[…]
In Vivo Pharmacokinetics of Evolving Hydrogels with Therapeutic Cargo in Mice
As an ultimate test of applicability in a translational setting, we used the surfactant-modified Sangelose hydrogels to obtain in vivo pharmacokinetics (PK) data (Fig. 5). We assessed the serum protein concentration over time following injection of Sangelose hydrogels loaded with hIgG into mice, and the data were analyzed using a two-compartment pharmacokinetics model to describe the release kinetics (details in the SI). Both groups exhibited typical pharmacokinetics profiles with an initial increase in serum protein concentration followed by a decline. S3Tw0 reached a peak concentration ($c_{\rm max}$) of $282.27\leavevmode\nobreak\ \mu{\rm g\leavevmode\nobreak\ mL}^{-1}$ at $48\leavevmode\nobreak\ {\rm h}$ ($t_{\rm max}$), while S3Tw2 reached a peak of $260.69\leavevmode\nobreak\ \mu{\rm g\leavevmode\nobreak\ mL}^{-1}$ at $96\leavevmode\nobreak\ {\rm h}$. It is clear that S3Tw0 demonstrated a faster release profile, with a more pronounced burst release, as indicated by the earlier $t_{\rm max}$ and higher $c_{\rm max}$. In contrast, S3Tw2 exhibited a more sustained release, with slower decline in serum concentrations, as evidenced by its larger area under the curve (AUC) of $55,181.69\leavevmode\nobreak\ \mu{\rm g\leavevmode\nobreak\ h\leavevmode\nobreak\ mL}^{-1}$, compared to $50,752.97\leavevmode\nobreak\ \mu{\rm g\leavevmode\nobreak\ h\leavevmode\nobreak\ mL}^{-1}$ for S3Tw0. Fits of the PK data (cf. Fig. 5(c)) to a two-compartment model also yield the half-lives for (i) depot/cargo release into the subcutaneous space and subsequent absorption into the bloodstream, $\tau_{\rm r}$ (compartment 1) and (ii) elimination from the body, $\tau_{\rm e}$ (compartment 2) (details in the SI). From the model fits, $\tau_{\rm r}$ was obtained to be $3.19\pm 1.27$ and $6.02\pm 1.04\leavevmode\nobreak\ {\rm d}$ for S3Tw0 and S3Tw2 respectively (cf. Fig. 5(e), whereas $\tau_{\rm e}$ was found to be $1.30\pm 0.40$ and $1.01\pm 0.16\leavevmode\nobreak\ {\rm d}$ respectively. Elimination kinetics are typically assumed to be independent of administration method and vehicle, and thus the values being essentially the same across both groups with very similar hydrogel structure, chemistry, and biological compatibility are consistent with previous literature (13, 35). The only significant difference between the two groups was observed in the release half-lives, where S3Tw2 released cargo significantly slower than S3Tw0. These findings highlight the improved burst-release performance of Sangelose functionalized with tween 80, achieving slower overall release while maintaining sustained cargo delivery.
Discussion
This study highlights the potential of injectable dynamic hydrogels to intrinsically evolve their transport properties post-injection within subcutaneous depots, offering a novel approach to mitigating burst release while modulating cargo release rates. By engineering the magnitude and direction of changes in rheological and mass transport properties through functional molecules that influence crosslinking kinetics, we achieve hybrid release profiles seamlessly transitioning between burst and sustained release. This self-regulated evolution addresses a critical challenge in injectable hydrogels, mitigating sudden, excessive cargo release while maintaining effective sustained delivery. Such doubly-controlled release profiles are beneficial for a diverse array of therapeutics deliveries: minimal burst release is strongly desirable for delivery of therapeutics with strong toxicities, e.g. GLP-1 agonists in diabetes treatment (36); slow, prolonged release is essential for ultra-long administration of precious drug cargo for treating various infectious diseases (13). Decoupling and independently controlling these release regimes overcomes the inherent constraints of stable crosslinked networks. Modular control over network relaxation, recovery, and solute diffusivity offers a blueprint for designing hydrogels tailored to the complex demands of drug delivery and other biomedical applications. Our findings establish a foundation for next-generation soft materials capable of dynamically adapting their structure and function to physiological conditions. This advancement enables more precise therapeutic delivery and bridges the gap between material design and clinical application. Future work can expand on these insights to explore broader applications and refine control mechanisms for complex release profiles, unlocking the full potential of injectable hydrogels in diverse biomedical settings.
Conflicts of Interest
E.A.A., S.S., C.D., and Y.E.S. are listed as inventors on patent applications describing the hydrogel technology reported in this manuscript. E.A.A. is a co-founder, equity holder, and advisor to Appel Sauce Studios LLC, which holds a global exclusive license to these technologies from Stanford University. All other authors declare no competing financial interests.
[…]
Materials
[…]
Sangelose hydrogels were prepared by dissolving Sangelose 90L (Daido Chemical Co., Osaka, Japan) powder in phosphate-buffered saline (PBS) to form a 6% w/w stock, which was then diluted to 3% w/w. Tween 80 was added during dilution as needed. Detailed formulation procedures are provided in the SI.
Shear Rheometry
Rheological measurements were conducted using a Discovery Hybrid Rheometer (DHR-2, TA Instruments). Details of rheometry procedures and continuous relaxation spectra analysis are provided in the SI.
In Vitro Diffusivity Measurements
The diffusion coefficients of probe cargoes in the hydrogels were determined using fluorescence recovery after photobleaching (FRAP), procedure is detailed in the SI. Fluorescein isothiocyanate-tagged bovine serum albumin (BSA, Sigma, $M_{\rm w}\simeq 68\leavevmode\nobreak\ {\rm kDa}$) was used for PVA-Borax, and Alexa 647-tagged human immunoglobulin (hIgG, Sigma, $M_{\rm w}\simeq 150\leavevmode\nobreak\ {\rm kDa}$) for Sangelose.
Briefly, fluorescent cargo was bleached with a high-intensity laser, and the fluorescence recovery time constant, $t_{1/2}$, was used to calculate the diffusion coefficient $\mathbb{D}=\gamma_{\mathbb{D}}(\rho^{2}/t_{1/2})$, where $\gamma_{\mathbb{D}}$ is an empirically determined constant (37).
In Vitro Release Assays
Release assays were conducted by injecting $100\leavevmode\nobreak\ \mu{\rm L}$ of PVA-Borax and Sangelose hydrogels loaded with FITC-BSA and Alexa-hIgG respectively into a capillary tube filled with $400\leavevmode\nobreak\ \mu{\rm L}$ of the appropriate buffer. At regular intervals, the supernatant was removed and analyzed to measure the amount of cargo present. Detailed procedures are described in the SI.
In Vivo Pharmacokinetics Studies
Animal studies were conducted on female BL6 mice (6-8 weeks old) administered hIgG via subcutaneous (SC) injection of 100 $\mu$L Sangelose hydrogel containing 1 mg hIgG/mouse. Serum samples were collected from the tail at 2 and 8 h, and 1, 2, 4, 7, 10, and 14 d ($n=4$) for both groups (S3Tw0 and S3Tw2), and analyzed for hIgG concentration by ELISA. Details of gel preparation, injection, ELISA, and pharmacokinetics modeling are provided in the SI. NIH guidelines for the care and use of laboratory animals (NIH Publication #85-23 Rev. 1985) have been observed. All animal studies were performed with the approval of Stanford Administrative Panel on Laboratory Animal Care.
Statistical Analyses
The statistical analyses to determine the significance levels across datasets for the in vitro and in vivo data were performed using a Kolmogorov-Smirnov test in MATLAB with a 5% threshold.
[…]
Supplementary Information and Datasets
All experimental data shown in the figures are provided as .xls files in the Supplementary Datasets section.
[…]
The authors thank the Bill & Melinda Gates Foundation (INV-008642) for their financial support of this work in the development of facile and effective injectable drug delivery platforms. The probe diffusivity data using FRAP were obtained using confocal microscopy at the Stanford Cell Sciences Imaging Facility (CSIF).
[…]
[40]